H. Cochran, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $3,100,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $3,370,000 in annual sales, with costs of $2,390,000. If the tax rate is 24 percent, what is the OCF for this project?
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(3,100,000/3)=$1,033,333.33

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(3,370,000-2,390,000)(1-0.24)+(0.24*1033333.33)

=$992800.